Determine the area of the region bounded by the given set of curves y=x√x^2 +1, y=e^-1/2, x=-3 and the y-axis.


1
Expert's answer
2021-04-15T07:34:06-0400

The region bounded by equation "y=x\u221a(x^2+1)" , "y=e^{-1\/2}" , "x=-3" and y-axis

"\therefore area= ∫^{0} _{-3} [(e) ^{-1/2} -x√(x^{2}+1)) ]dx"

"=\u222b_{-3}^{0}e^{-1\/2}dx- \u222b_{-3}^{0}x\u221a(x^2+1 ) dx"

"=[ e^{-1\/2} x]|_{-3}^{0}-1\/2 \u222b_{-10}^{1} u^{1\/2} du"

letting "u=x^2+1" in the second integral and substituting up to limits 

"=[0-e^{-1/2} (-3)]- \frac{1}{2} [\frac{2}{3}u^{3/2} )]|^{1} _{10}"

"=3e^{-1/2}- \frac{1}{3} [1^{3/2 } -10^{3/2} ]"

"=3e^{-1/2}+ \frac{10}{3} √10-\frac{1}{3}"

 = 12.03 square units of length